5. Ravi is trying to find out which investment is better. Investment option 1 states interest at 11.10% p.a. compounded quarterly. Investment option 2 states interest at 10.90% p a. compounded monthly. Which is the better investment choice?


1
Expert's answer
2021-04-19T18:53:02-0400
"(1+\frac{0.111}{12})^4=1.0373"

"(1+\frac{0.109}{12})^{12}=1.115"

Need to choice second progect.